OpenStudy (anonymous):

Solve. Problem in comments. A. s ≥ –2 B. s ≤ –2 C. s ≥ –18 D. s ≤ –18

OpenStudy (anonymous):

\[-\frac{ s }{ 3 } ≥ 6\]

OpenStudy (anonymous):

multiply both side by 3: \[-s \ge 18\] multiply both sides by -1. Important: when you do this you must turn the sign around too: \[s \le -18\] to confirm your answer. plug in random s less than or equal to -18 in the initial equation to make sure it holds true

OpenStudy (anonymous):

thanks :)